8) A wine-dispensing system uses argon canisters to pressurize and preserve wine in the bottle. An argon canister for the system has a volume of 60.1 mL and contains 21.2 g of argon. Assuming ideal gas behavior, what is the pressure in the canister at 317 K ? Express your answer with the appropriate units.

B) When the argon is released from the canister it expands to fill the wine bottle. How many 750.0-mL wine bottles can be purged with the argon in the canister at a pressure of 1.32 atm and a temperature of 317 K ?
Answer #1

A) According to ideal gas equation:

PV = (m/M)RT

Where 'M' is the molar mass of argon = 40 g/mol

i.e. P * 60.1*10-3 L = 21.2 g/(40 g/mol) * 0.08206 L.atm/mol.K * 317 K

Therefore, the pressure in the canister (P) = 229.4 atm

B) According to ideal gas equation: PV = nRT

1.32 atm * 750*10-3 L = n * 0.08206 L.atm/mol.K * 317 K

Therefore, the no. of wine bottles that can be purged with argon (n) = 0.038 mol = 0.038 * 6.022*1023 = 2.29*1022